A force of 80 newton is working between two point charges placed at a distance in air when these charges are placed at the same distance in a dielectric medium then the force become 8 newton Find the dielectri constant of the medium

Grade:12

1 Answers

Khimraj
6 years ago
Let the dielectric constant of dielectric medium is K. (K = ε/ε0)
Force between two point charges placed at a distance in air is F = 80N
Force when these placed at same distance in dielectric medium = F’ = F/K = 8N
so (80N)/K = 8N
then K =10.
Hope it clears. If still  you have doubt you can ask me. If you like answer then please approve it.